### v2

@@ 2 Model
This paper considers mechanisms that produce matching functions with a cutoff characterization according to Definition 2. Placement scores $\bm{\mathrm{S}}$ may or may not equal school priority scores $\bm{\mathrm{R}}$. The definition gives researchers the freedom to construct special placement scores $\bm{\mathrm{S}}$ if the mechanism that they consider does not admit cutoff characterization by means of priority scores $\bm{\mathrm{R}}$. The idea behind this definition stems from the logic of the general class of mechanisms of agarwal2018. Moreover, we assume throughout the paper that, for any $j\in\mathcal{J}$, $S_{j}$ has a distribution that is absolutely continuous with respect to (wrt) the Lebesgue measure and support $\mathcal{S}_{j}$ that contains a closed interval around $c_{j}$. The support of $\bm{\mathrm{S}}$ is $\mathcal{S}$. For any two scores $S_{j}$ and $S_{l}$, we assume that either $\mathbb{P}[S_{j}=S_{l}]=1$ or $\mathbb{P}[f(S_{j})=S_{l}]<1$ for any measurable function $f$. This says that the only deterministic function relating any two scores may be the identity function.
azevedo2016 demonstrate that, if the matching function is stable, the matching has cutoff characterization with $\bm{\mathrm{S}}=\bm{\mathrm{R}}$ and admission cutoffs constructed as follows. For each $j\in\mathcal{J}$, $c_{j}:=\inf\{S_{j}(\omega):\text{ for }\omega\text{ with }\mu(\omega)=j\}$ if some individuals are matched to $j$ or $c_{j}:=\inf\mathcal{S}_{j}$ if nobody is matched to $j$. Many mechanisms produce stable matchings. For example, SD and DA are strategy proof and lead to stable matchings if agents are truth-tellers. Regarding settings where agents are not truth-tellers, e.g., because they face constraints in the submission of $P$, haeringer2009 study how the SD, DA, Boston, and TTC mechanisms lead to stable matchings.
@@
This paper is about identification of moments of treatment effects $Y(d^{\prime})-Y(d)$. The econometrician has access to an infinite amount of data and observes the joint distribution of the following random objects: $P(\omega)$, $\bm{\mathrm{S}}(\omega)$, $\mu(\omega)$, and $Y(\omega):=Y(\omega,\mu(\omega))$, where we abuse the notation and employ the letter $Y$ for both the observed outcome, $Y(\omega)$, and the potential outcome of being assigned to school $d$, $Y(\omega,d)$.
@@ 3 Identification with Truthful Reports
In this section, we consider identification of causal effects when all students are truth-tellers, that is, when they submit their true ranking of acceptable schools. This behavior is rational in strategy-proof mechanisms. For instance, dubins1981 and roth1982 show that the DA mechanism without constraints on preference submission is strategy proof in the finite economy; abdul2015 show the same for the continuum economy.555 See also the work by azevedo2019, who advocate a robust notion of strategyproofness in a large economy. The SD is also strategy proof, as it is a special case of DA.
@@
Students’ submitted list of acceptable schools reflects their true preferences.
Although submitting the true list of acceptable schools is a weakly dominant strategy under strategy-proof mechanisms, Assumption 1 explicitly contrasts with recent evidence indicating that students do otherwise (chen2006, pais2008).
@@
Unlike in a standard sharp RD, that $S_{j}(\omega)$ clears the cutoff $c_{j}$ does not automatically determine that student $\omega$ is allocated to school $j$. This is the case only when $j$ is the most preferred school among the schools that the student prefers, that is, when $j$ is the favorite school in the set of schools for which the student clears the cutoff.
@@
The first step in the RD is to correctly identify the marginal individuals for a given cutoff and a given change in school, e.g., cutoff $c_{j}$ and a change from school $k$ to school $j$. For example, for any individual with score $S_{j}$ just to the right of $c_{j}$, we need to determine two things: that the individual is matched to school $j$, and that the individual would have been matched to school $k$ had her score been just to the left of $c_{j}$. The cutoff representation implies that these two things depend on the counterfactual sets of available schools on either side of the cutoff and on the individual’s preferences over these sets.
@@
It is straightforward to obtain counterfactual sets of available schools in the case where all schools rely on the same placement score, that is, $S_{j}=S_{1}$ for every $j$. For example, this is the case under the SD mechanism. In this case, the set of feasible schools is all schools with a cutoff below or equal to score $S_{1}$. Note that everyone just above (or just below) cutoff $c_{j}$ has exactly the same set of feasible schools. For someone with $S_{1}\geq c_{j}$, the counterfactual scenario has the score crossing to the left of cutoff $c_{j}$, and school $j$ is dropped from the set of feasible schools. In turn, for someone with $S_{1}<c_{j}$, the counterfactual scenario adds school $j$ to the set of feasible schools.666 A common practice in applied work consists of “cleaning” irrelevant schools from the submitted preference lists in cases where $S_{j}=S_{1}$ for every $j$. For instance, say an individual submits $P=\{1,2,3\}$ and $c_{2}>S_{1}>c_{1}>c_{3}$. Given the cutoff characterization and truth-telling, the matching assignment of this individual is school $1$; the counterfactual assignment is school $3$ even though $32P$ 3; this is the case because school $2$ has a cutoff higher than the cutoff of school $1$. In this case, the irrelevant school to be cleaned from $P$ is school $2$. The general idea is to remove all schools ranked below $1$ that have cutoffs higher than $c_{1}$. See the description of this practice by estrada2017. The practice cannot be used to identify counterfactual assignments in cases where different schools use different placement scores, as under, e.g., the DA mechanism. Unlike under SD, agents near and on the same side of a cutoff in DA differ in their sets of feasible schools. It is not immediately obvious which schools appear in their counterfactual sets. In DA, schools use different scores, and these scores may be functions (e.g., weighted averages) of a small set of primitive scores. That is the case in our application with the Chilean data. This makes the joint support of the distribution of scores highly dependent and complicates the counterfactual analysis. Dealing with this complexity is empirically relevant since many real-world higher education assignment mechanisms use DA. See Table 1 Panel B in fack2019 for a list of examples.
@@
Example (SD Example).
In SD, $S_{j}=S_{1}$ for every $j$, and the definition of the budget set above equals $B(\bm{\mathrm{S}})=\{m:S_{1}\geq c_{m}\}$. Suppose we have four schools with cutoffs $c_{1}<c_{2}<c_{3}<c_{4}$. Individuals in this economy have five possible budget sets: $\{0\}$, $\{0,1\}$, $\{0,1,2\}$, $\{0,1,2,3\}$, and $\{0,1,2,3,4\}$. For individuals near cutoff $c_{4}$, the counterfactual budget sets are $B^{-}_{4}(\bm{\mathrm{S}})=\{0,1,2,3\}$ and $B^{+}_{4}(\bm{\mathrm{S}})=\{0,1,2,3,4\}$.777 In SD or in DA with independent placement scores, the definitions of the counterfactual equal $B^{+}_{j}(\bm{\mathrm{S}})=B(\bm{\mathrm{S}})\cup\{m:c_{m}=c_{j}\}$ and $B^{-}_{j}(\bm{\mathrm{S}})=B(\bm{\mathrm{S}})\setminus\{m:c_{m}=c_{j}\}$; moreover, if the cutoffs are unique, $B^{+}_{j}(\bm{\mathrm{S}})=B(\bm{\mathrm{S}})\cup\{j\}$ and $B^{-}_{j}(\bm{\mathrm{S}})=B(\bm{\mathrm{S}})\setminus\{j\}$.
Next, we define the concept of local preferences, that is, the first- and second-best choices for a marginal individual at any given cutoff.
@@
Example (SD Example, Continued).
Consider four individuals whose submitted preferences are $P^{(1)}=\{3,4,2,1\}$, $P^{(2)}=\{4,1,2,3\}$, $P^{(3)}=\{4,2,3,1\}$, and $P^{(4)}=\{4,3,1,2\}$. Their corresponding local preferences are $P^{(1)}_{4}=(3,3)$, $P^{(2)}_{4}=(4,1)$, $P^{(3)}_{4}=(4,2)$, and $P^{(4)}_{4}=(4,3)$.
There is no distinction between $P_{j}$ and $Q_{j}$ at this stage because of Assumption 1. Students are truthful when they submit their list of acceptable schools, so $P$ equals the schools listed higher than 00 in $Q$ and $P_{j}=Q_{j}$. Section 4 considers the case of strategic misreporting. In that case, $P_{j}$ is observed but $Q_{j}$ is not. Cutoff characterization implies that an individual with $Q_{j}=(k,l)$ is matched to school $k$ if $S_{j}$ is just above $c_{j}$ or to school $l$ if $S_{j}$ is just below $c_{j}$. The same applies for individuals with $P_{j}=(k,l)$ under Assumption 1.
@@
We adopt the convention that comparable pairs do not involve the outside option 00. We do this because it may be hard to interpret the treatment effects of a change from the outside option to a school when the outside option varies across individuals. We do not consider pairs with $j=k$ because the initial school assignment does not change for these individuals. We also exclude pairs $Q_{j}=(j^{\prime},k)$ with $j^{\prime}\neq j$ from $\mathcal{P}$ to avoid redundancy. We may find individuals with $Q_{j}=(j^{\prime},k)$ whenever schools $j$ and $j^{\prime}$ use the same score and have the same cutoff. As the score $S_{j}=S_{j^{\prime}}$ crosses the cutoff $c_{j}=c_{j^{\prime}}$, access is granted to both schools $j$ and $j^{\prime}$, and individuals may differ in their preferences for these schools. Individuals who prefer $j^{\prime}$ will not appear in $\mathcal{P}$ as having $Q_{j}=(j^{\prime},k)$, but they may appear in $\mathcal{P}$ with $Q_{j^{\prime}}=(j^{\prime},k)$.
The purpose of defining counterfactual sets and local preferences is to construct a local preference variable for every student and use it as a control variable in the RD. In this section, this variable is $P_{j}$, which equals to $Q_{j}$ because of Assumption 1. When we focus on students with $P_{j}=(j,k)$, $k\neq j$, the marginal switch in the allocation around cutoff $c_{j}$ becomes a function of $S_{j}$. Controlling for $P_{j}$ ensures that we apply the RD strategy to all individuals whose assignment switches from $k$ to $j$ at the cutoff; this identifies the effect of the change in the assignment as long as the typical RD continuity assumptions are satisfied.
RD identification requires continuity assumptions on the distribution of individual types conditional on the relevant placement score. In our case, we also need to verify continuity after we condition on $P_{j}=Q_{j}$. After all, we do not want to condition on a characteristic inconsistent with the central argument for identification in RD: that individuals to the right and the left of the cutoff are “similar on average”. Below, we state an assumption on the continuity of types and prove that it implies the kind of smoothness required by RD. Before we do so, we define the following set of events. For every school $j\in\mathcal{J}$, partition the placement scores $\bm{\mathrm{S}}$ as the score of school $j$ and all other scores: $\bm{\mathrm{S}}\equiv(S_{j},\bm{\mathrm{S}}_{-j})$. Define $\bm{\mathrm{A}}_{-j}$ to be the collection of events on $\bm{\mathrm{S}}_{-j}$ that determine the availability of all non-$j$ schools. There are $2^{J-1}$ such events in $\bm{\mathrm{A}}_{-j}$. For example, if $J=2$, $\bm{\mathrm{A}}_{-1}=\{\{S_{2}\geq c_{2}\},\{S_{2}<c_{2}\}\}$; if $J=3$, $\bm{\mathrm{A}}_{-1}=\{\{S_{2}\geq c_{2},S_{3}\geq c_{3}\},\{S_{2}\geq c_{2},S_{3}<c_{3}\},\{S_{2}<c_{2},S_{3}\geq c_{3}\},\{S_{2}<c_{2},S_{3}<c_{3}\}\}$; and so forth.
@@
Proposition 1 shows that a standard RD is valid in the truth-telling case as long as we control for $P_{j}$. The parameter of interest is the average treatment effect on $g(Y)$ from a change in the school of assignment from $j$ to $k$, averaged over individuals at the cutoff $c_{j}$ and with true local preferences $(j,k)$. Indeed, these are the parameters of interest for the rest of this paper:
@@
Parameters like these are of economic interest not only for evaluating current mechanisms in place but also for studying the effects of counterfactual assignments (e.g., of students near $c_{j}$ currently at school $k$ being offered admission at school $j$.) The fact that the parameter is conditional on $Q_{j}$ instead of $P_{j}$ is important for the external validity of (1) when $P$ is manipulable by agents.

### v3

@@ 2 Model
This paper considers mechanisms that produce matching functions with a cutoff characterization according to Definition 2. Placement scores $\bm{\mathrm{S}}$ may or may not equal school priority scores $\bm{\mathrm{R}}$. The definition gives researchers the freedom to construct special placement scores $\bm{\mathrm{S}}$ if the mechanism that they consider does not admit cutoff characterization by means of priority scores $\bm{\mathrm{R}}$. The idea behind this definition stems from the logic of the general class of mechanisms of Agarwal and Somaini (2018). Moreover, we assume throughout the paper that, for any $j\in\mathcal{J}$, $S_{j}$ has a distribution that is absolutely continuous with respect to (wrt) the Lebesgue measure and support $\mathcal{S}_{j}$ that contains a closed interval around $c_{j}$. The support of $\bm{\mathrm{S}}$ is $\mathcal{S}$. For any two scores $S_{j}$ and $S_{l}$, we assume that either $\mathbb{P}[S_{j}=S_{l}]=1$ or $\mathbb{P}[f(S_{j})=S_{l}]<1$ for any measurable function $f$. This says that the only deterministic function relating any two scores may be the identity function. Azevedo and Leshno (2016) demonstrate that, if the matching function is stable, the matching has cutoff characterization with $\bm{\mathrm{S}}=\bm{\mathrm{R}}$ and admission cutoffs constructed as follows. For each $j\in\mathcal{J}$, $c_{j}:=\inf\{S_{j}(\omega):\text{ for }\omega\text{ with }\mu(\omega)=j\}$ if some individuals are matched to $j$ or $c_{j}:=\inf\mathcal{S}_{j}$ if nobody is matched to $j$. Many mechanisms produce stable matchings. For example, SD and DA are strategy proof and lead to stable matchings if agents are truth-tellers. Regarding settings where agents are not truth-tellers, e.g., because they face constraints in the submission of $P$, Haeringer and Klijn (2009) study how the SD, DA, Boston, and TTC mechanisms lead to stable matchings.
Parameter of Interest
Having laid out the physical environment of our matching economy, we now motivate the type of causal parameter that concerns this paper. Causal parameters are often used as policy tools for their internal validity—evaluating the impact of historical interventions on outcomes—or for their external validity—assessing the impact of new interventions based on knowledge of historical interventions, that is, constructing counterfactual analyses; see Heckman (2005) for a detailed discussion.
@@
In this paper, we aim to provide strategies for identifying a set of parameters that are not only internally valid but also possess substantial external validity, thereby enabling the evaluation of counterfactual policies. We seek to identify conditional moments of treatment effects $Y(d^{\prime})-Y(d)$ when the econometrician has access to an infinite amount of data and observes the joint distribution of the following random objects: $P(\omega)$, $\bm{\mathrm{S}}(\omega)$, $\mu(\omega)$, and $Y(\omega):=Y(\omega,\mu(\omega))$.555 We abuse the notation and employ the letter $Y$ for both the observed outcome, $Y(\omega)$, and the potential outcome of being assigned to school $d$, $Y(\omega,d)$.
Section A.1 in the appendix provides a detailed discussion of a class of counterfactual analyses that is advocated by Artemov et al. (2023). Analyzing the welfare effects of a policy change in that class relies on identification of the average structural functions (ASF)
$s\mapsto\mathbb{E}[Y(j)\mid Q=q,\;S_{j}=s]$ (1)
for $j\in\{0,\ldots,J\}$ and $q\in\mathcal{Q}$. Nonparametric identification of these functions is extremely challenging. However, this paper shows that combining an RD identification strategy with data from assignment mechanisms that satisfy our assumptions yields set identification of differences of the ASFs in (1); in other words, our proposed methods set-identify
$\mathbb{E}[Y(j)-Y(k)\mid Q=q,\;S_{j}=c_{j}]$ (2)
at finitely many values of $j$, $k$, and $q$ (see Propositions 1 and 5).
The parameters in (2) are the object of interest of this paper. They carry the simple intuition of the massively popular RD designs while providing useful identifying information for a large class of counterfactual policy evaluations. Section A.1 in the appendix describes how researchers may combine all information identified by RD with smoothness assumptions on the ASFs to construct bounds on a variety of counterfactual policy effects.
@@ 3 Identification with Truthful Reports
In this section, we consider identification of causal effects when all students are truth-tellers, that is, when they submit their true ranking of acceptable schools. We start with truth-telling in order to introduce our approach and notation in a simple behavioral setting before moving to strategic reports in Section 4. Truth-telling is rational in strategy-proof mechanisms. For instance, Dubins and Freedman (1981) and Roth (1982) show that the DA mechanism without constraints on preference submission is strategy proof in the finite economy; Abdulkadiroglu et al. (2015) show the same for the continuum economy.666 See also the work by Azevedo and Budish (2019), who advocate a robust notion of strategyproofness in a large economy. The SD is also strategy proof, as it is a special case of DA. On the other hand, experimental evidence show that some individuals deviate from the rational behavior of truth-telling (Chen and Sönmez (2006), Pais and Pintér (2008)). Thus, we explicitly assume everyone is a truth-teller.
@@
Students submit their true list of acceptable schools.
@@
Unlike in a standard sharp RD, that $S_{j}(\omega)$ clears the cutoff $c_{j}$ does not automatically determine that student $\omega$ is allocated to school $j$. This is the case only when $j$ is the most preferred school among the schools that are feasible to the student, that is, when $j$ is the favorite school in the set of schools for which the student clears the cutoff.
@@
The first step in the RD is to correctly identify the marginal individuals for a given cutoff and a given change in schools. For example, for any individual with score $S_{j}$ just to the right of $c_{j}$, we need to determine two things: that the individual is matched to school $j$, and that the individual would have been matched to school $k$ had her score been just to the left of $c_{j}$. The cutoff representation implies that these two things depend on the counterfactual sets of available schools on either side of the cutoff and on the individual’s preferences over these sets.
@@
It is straightforward to obtain counterfactual sets of available schools in the case where all schools rely on the same placement score, that is, $S_{j}=S_{1}$ for every $j$. For example, this is the case under the SD mechanism. In this case, the set of feasible schools is all schools with a cutoff below or equal to score $S_{1}$. Note that everyone just above (or just below) cutoff $c_{j}$ has exactly the same set of feasible schools. For someone with $S_{1}\geq c_{j}$, the counterfactual scenario has the score crossing to the left of cutoff $c_{j}$, and school $j$ is dropped from the set of feasible schools. In turn, for someone with $S_{1}<c_{j}$, the counterfactual scenario adds school $j$ to the set of feasible schools. Unlike under SD, agents near and on the same side of a cutoff in DA differ in their sets of feasible schools. It is not immediately obvious which schools appear in their counterfactual sets. In DA, schools use different scores, and these scores may be functions (e.g., weighted averages) of a small set of primitive scores. That is the case in our application with the Chilean data. This makes the joint support of the distribution of scores highly dependent and complicates the counterfactual analysis. Dealing with this complexity is empirically relevant since many real-world higher education assignment mechanisms use DA. See Table 1 Panel B in Fack et al. (2019) for a list of examples.777 A common practice in applied work consists of “cleaning” irrelevant schools from the submitted preference lists in cases where $S_{j}=S_{1}$ for every $j$. For instance, say an individual submits $P=\{1,2,3\}$ and $c_{2}>S_{1}>c_{1}>c_{3}$. Given the cutoff characterization and truth-telling, the matching assignment of this individual is school $1$; the counterfactual assignment when $c_{2}>c_{1}>S_{1}>c_{3}$ is school $3$ even though $2P3$; this is the case because school $2$ has a cutoff higher than the cutoff of school $1$. In this case, the irrelevant school to be cleaned from $P$ is school $2$. The general idea is to remove all schools ranked below $1$ that have cutoffs higher than $c_{1}$. See the description of this practice by Estrada and Gignoux (2017). The practice cannot be used to identify counterfactual assignments in cases where different schools use different placement scores, as under, e.g., the DA mechanism.
@@
Example (SD Example, Part I).
In SD, $S_{j}=S_{1}$ for every $j$, and the definition of the budget set above equals $B(\bm{\mathrm{S}})=\{m:S_{1}\geq c_{m}\}$. Suppose we have four schools with cutoffs $c_{1}<c_{2}<c_{3}<c_{4}$. Individuals in this economy have five possible budget sets: $\{0\}$, $\{0,1\}$, $\{0,1,2\}$, $\{0,1,2,3\}$, and $\{0,1,2,3,4\}$. For individuals near cutoff $c_{4}$, the counterfactual budget sets are $B^{-}_{4}(\bm{\mathrm{S}})=\{0,1,2,3\}$ and $B^{+}_{4}(\bm{\mathrm{S}})=\{0,1,2,3,4\}$.888 In SD or in DA with independent placement scores, the definitions of the counterfactual equal $B^{+}_{j}(\bm{\mathrm{S}})=B(\bm{\mathrm{S}})\cup\{m:c_{m}=c_{j}\}$ and $B^{-}_{j}(\bm{\mathrm{S}})=B(\bm{\mathrm{S}})\setminus\{m:c_{m}=c_{j}\}$; moreover, if the cutoffs are unique, $B^{+}_{j}(\bm{\mathrm{S}})=B(\bm{\mathrm{S}})\cup\{j\}$ and $B^{-}_{j}(\bm{\mathrm{S}})=B(\bm{\mathrm{S}})\setminus\{j\}$.
Next, we follow the intuition of Kirkeboen et al. (2016) and define the concept of local preferences, that is, the first- and second-best choices for a marginal individual at any given cutoff. This will later become the control variable in our RD identification strategy with truthful agents.
@@
Example (SD Example, Part II).
Consider four individuals whose submitted preferences are $P^{(1)}=\{3,4,2,1\}$, $P^{(2)}=\{4,1,2,3\}$, $P^{(3)}=\{4,2,3,1\}$, and $P^{(4)}=\{4,3,1,2\}$. Their corresponding local preferences at cutoff $c_{4}$ are $P^{(1)}_{4}=(3,3)$, $P^{(2)}_{4}=(4,1)$, $P^{(3)}_{4}=(4,2)$, and $P^{(4)}_{4}=(4,3)$.
There is no distinction between $P_{j}$ and $Q_{j}$ at this stage because of Assumption 1. Students are truthful when they submit their list of acceptable schools, so $P$ equals the schools listed higher than 0 in $Q$ and $P_{j}=Q_{j}$. Section 4 considers the case of strategic misreporting. In that case, $P_{j}$ is observed but $Q_{j}$ is not. Cutoff characterization implies that an individual with $Q_{j}=(k,l)$ is matched to school $k$ if $S_{j}$ is just above $c_{j}$ or to school $l$ if $S_{j}$ is just below $c_{j}$. The same applies for individuals with $P_{j}=(k,l)$ under Assumption 1.
@@
We adopt the convention that comparable pairs do not involve the outside option 0. We do this because it may be hard to interpret the treatment effects of a change from the outside option to a school when the outside option varies across individuals. We do not consider pairs with $j=k$ because the initial school assignment does not change for these individuals. We also exclude pairs $Q_{j}=(j^{\prime},k)$ with $j^{\prime}\neq j$ from $\mathcal{P}$ to avoid redundancy. We may find individuals with $Q_{j}=(j^{\prime},k)$ whenever schools $j$ and $j^{\prime}$ use the same score and have the same cutoff. As the score $S_{j}=S_{j^{\prime}}$ crosses the cutoff $c_{j}=c_{j^{\prime}}$, access is granted to both schools $j$ and $j^{\prime}$, and individuals may differ in their preferences for these schools. Individuals who prefer $j^{\prime}$ will not appear in $\mathcal{P}$ as having $Q_{j}=(j^{\prime},k)$, but they may appear in $\mathcal{P}$ with $Q_{j^{\prime}}=(j^{\prime},k)$.
The purpose of defining counterfactual sets and local preferences is to construct a variable for every student and use it as a control variable in the RD. In this section, this variable is $P_{j}$, which equals to $Q_{j}$ because of Assumption 1. When we focus on students with $P_{j}=(j,k)$, $k\neq j$, the marginal switch in the allocation around cutoff $c_{j}$ becomes a function of $S_{j}$. Controlling for $P_{j}$ ensures that we apply the RD strategy to all individuals whose assignment switches from $k$ to $j$ at the cutoff; this identifies the effect of the change in the assignment as long as the typical RD continuity assumptions are satisfied.
RD identification requires continuity assumptions on the distribution of individual types conditional on the relevant placement score. In our case, we also need to verify continuity after we condition on $P_{j}=Q_{j}$. After all, we do not want to condition on a variable that breaks the central argument for identification in RD: that individuals to the right and the left of the cutoff are “similar on average”. Below, we state an assumption on the continuity of types and prove that it implies the kind of smoothness required by RD. Before we do so, we define the following set of events. For every school $j\in\mathcal{J}$, partition the placement scores $\bm{\mathrm{S}}$ as the score of school $j$ and all other scores: $\bm{\mathrm{S}}\equiv(S_{j},\bm{\mathrm{S}}_{-j})$. Define $\bm{\mathrm{A}}_{-j}$ to be the collection of events on $\bm{\mathrm{S}}_{-j}$ that determine the availability of all non-$j$ schools. There are $2^{J-1}$ such events in $\bm{\mathrm{A}}_{-j}$. For example, if $J=2$, $\bm{\mathrm{A}}_{-1}=\{\{S_{2}\geq c_{2}\},\{S_{2}<c_{2}\}\}$; if $J=3$, $\bm{\mathrm{A}}_{-1}=\{\{S_{2}\geq c_{2},S_{3}\geq c_{3}\},\{S_{2}\geq c_{2},S_{3}<c_{3}\},\{S_{2}<c_{2},S_{3}\geq c_{3}\},\{S_{2}<c_{2},S_{3}<c_{3}\}\}$; etc.
@@
Proposition 1 shows that a standard RD is valid in the truth-telling case as long as we control for $P_{j}$. The parameter of interest is the average treatment effect on $g(Y)$ from a change in the school of assignment from $j$ to $k$, averaged over individuals at the cutoff $c_{j}$ and with true local preferences $(j,k)$ —fomally:
@@
Note that our parameter of interest (3) does not condition on the full preference profile  $Q=q$ as in (2) but on local preferences $Q_{j}=(j,k)$. We do so having in mind the data constraints inherent to the local nature of RD estimation. It follows that one value of $Q_{j}$ maps to multiple values of $Q$ for individuals with $S_{j}=c_{j}$ and (3) equals weighted averages of (2) over values of $Q$ (see proof of Lemma 1 in Section B.1 of the appendix).